OpenStudy (osgood):

Will medal!!!!! please help me with these questions Simplify the expressions 6. -4-5(x+8) -5x-44***I think it is this one -5x+36 -x-8 -20x+8 7. 2x+3(x-2)-3(x-6) 2x+12 2x-8 8x-24 2x-24 Thank you :)

OpenStudy (solomonzelman):

Yes, \(\text{#}6\) is correct.

OpenStudy (solomonzelman):

What do you think about \(\text{#}7\) ?

OpenStudy (osgood):

C? 8x-24?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 2x+3(x-2)-3(x-6) }\) \(\color{#000000 }{ \displaystyle 2x+3x-6-3x-(-18) }\)

OpenStudy (osgood):

?

OpenStudy (solomonzelman):

simplify that, and tell me what you get after that...

OpenStudy (osgood):

I am not good at simplifying things with parenthesis

OpenStudy (osgood):

That is why I asked for help :\

OpenStudy (solomonzelman):

Well, just \(\color{#000000 }{ \displaystyle -a(b-c)=-ab+ac }\) And \(\color{#000000 }{ \displaystyle a(b-c)=ab-ac }\)

OpenStudy (osgood):

@DariusX can you help me

OpenStudy (dariusx):